OpenStudy (anonymous):

2/5 (x − 4) = 2x for x

OpenStudy (anonymous):

i got 1/9 but its not one of my choices

OpenStudy (campbell_st):

multiply both sides by 5 2(x -4) = 10x expand the brackets 2x - 8 = 10x subtract 2x from both sides -8 = 8x divide both sides by 8 x = -1

jhonyy9 (jhonyy9):

cross multiplie 2(x-4)=10x 2x-8=10x -8=8x x=-1

OpenStudy (anonymous):

Thanks i totally did that wrong i turned everything into a fraction :(

OpenStudy (anonymous):

\[2(x-4)/5 = 2x\]\[2x-8=10x\]\[-8x=-8\]\[x=-1\]